OpenStudy (anonymous):

L'Hospital's Rule lim_x->(pi/2)^(-)_cosx*sec(5x)

OpenStudy (anonymous):

i got 1/5 but cramster says its 25...whos right?

OpenStudy (anonymous):

is the equation f(x)=-cos(x)*sec(5x)?

OpenStudy (anonymous):

who's* (for you grammar nazis)

OpenStudy (anonymous):

oh sorry forgot to elaborate... the equation is just cos(x)*sec(5x)

OpenStudy (anonymous):

the negative means its coming from the left side

OpenStudy (anonymous):

0/0 - differentiate both till you can eliminate the 0/0 condition

OpenStudy (anonymous):

oh no... the negative is actually a constant. you can factor the negative out when you do the limit thing. As such\[(-1)\lim_{x \rightarrow \pi/2}\cos(x)*\sec(5x)\]

OpenStudy (anonymous):

not its suppose to be \[\lim_{x \rightarrow (\pi/2)^{-}}\cos(x)*\sec(5x)\]

OpenStudy (anonymous):

no negatives...sorry i'm new here. and I don't see option to insert equation when u first type a question lol

OpenStudy (anonymous):

but i don't think "coming from the left" matters in this situation right??

OpenStudy (anonymous):

yup... it doesn't matter actually..... And yours answer is actually right...

OpenStudy (anonymous):

thanks!! thats all i needed to know!

OpenStudy (anonymous):

1/5 reconfirmation

OpenStudy (anonymous):

thanks :) looks like we have to correct cramster on this one